# Klear: Unified Multi-Task Audio-Video Joint Generation]Kling Team, Kuaishou Technology

### v1

@@ Abstract
Audio–video joint generation has progressed rapidly, yet substantial challenges still remain. Non-commercial approaches still suffer audio-visual asynchrony, poor lip–speech alignment, and unimodal degradation, which can be stemmed from weak audio–visual correspondence modeling, limited generalization, and scarce high-quality dense-caption data. To address these issues, we introduce Klear and delve into three axes—model architecture, training strategy, and data curation. Architecturally, we adopt a single-tower design with unified DiT blocks and an Omni-Full Attention mechanism, achieving tight audio–visual alignment and strong scalability. Training-wise, we adopt a progressive multitask regime—random modality masking to joint optimization across tasks, and a multistage curriculum, yielding robust representations, strengthening A-V aligned world knowledge, and preventing unimodal collapse. For datasets, we present the first large-scale audio–video dataset with dense captions, and introduce a novel automated data-construction pipeline which annotates and filters millions of diverse, high-quality, strictly aligned audio–video–caption triplets. Building on this, Klear scales to large datasets, delivering high-fidelity, semantically and temporally aligned, instruction-following generation in both joint and unimodal settings while generalizing robustly to out-of-distribution scenarios. Across tasks, it substantially outperforms prior methods by a large margin and achieves performance comparable to Veo 3, offering a unified, scalable path toward next-generation audio–video synthesis.
@@ 1 Introduction
The Kinetoscope was the world’s first film projector, invented by Thomas Edison in 1891. The Kinetophone, on the other hand, was the first sound film projector, combining the Kinetoscope with synchronized sound playback technology, marking a major breakthrough in sound film technology. Accordingly, vision and auditory are complementary and jointly indispensable modalities for depicting the real world. Consequently, audio–video joint generation has emerged as one of the most prominent trends in generative AI. Commercial systems such as Veo 3 and Sora 2 have achieved impressive performance with strong semantic alignment, while diverse open-source models with varying architectures are also emerging. Nevertheless, research on joint audio–visual generation remains nascent, and current models—even some commercial systems—still exhibit audio–visual asynchrony, lip–speech mismatches, and degradation in unimodal quality. We attribute several factors: (1) Architecture. Most T2AV models like JavisDiT [1], UniVerse-1 [2] and Ovi [3] employ a single-tower architecture with a cross-attention module, resulting in limited audio-video interaction and alignment. (2) Training strategies and data. Currently, mainstream methods perform single-task training, which might bring about biased representations and struggle to exploit underlying audio-video correlation and world knowledge.
@@
To address these issues, we propose Klear, which introduces coordinated improvements at the architectural, learning-strategy, and data levels. Specifically, on the architectural side, we adopt a single-tower backbone with unified DiT blocks. Each block integrates an Omni-Full Attention module that jointly attends to four streams—audio, audio captions, video, and video captions—thereby facilitating cross-modal fusion and interaction, achieving tight A-V alignment and stronger coupling to textual conditions, which offers a higher scaling ceiling. For training, we design a progressive multitask training regime: random modality masking sustains joint optimization over T2AV/TI2AV/TI2V/T2V/T2A. A performance-adaptive pretrain–post train curriculum tunes data mixtures and quality, yielding robust, generalizable representations that exploit A/V correlations and world knowledge. From the data construction, we introduce an automated annotation pipeline that enables efficient model scaling. Building on these three insights, our model achieves strong performance in both joint and unimodal generation, delivering high fidelity, tight audio–visual alignment, natural outputs, and robust instruction following with favorable scaling behavior. On the Verse-Bench, it surpasses prior methods by a large margin and generalizes well to out-of-distribution (OOD) scenarios. Our main contributions are summarized as follows:
  • We introduce Klear, a unified framework for multi-task audio–video joint generation that, to our knowledge, is the first model to achieve performance comparable to Veo 3. Klear effectively resolves semantic and temporal audio–visual misalignment while delivering high-fidelity generation.
@@
  • Extensive experiments demonstrate that Klear consistently excels in both unimodal and joint audio–video generation, outperforming prior state-of-the-art on the unimodal benchmark and the AV joint-generation benchmark consistently.
@@ 2 Related Works
Text-to-Video Generation (T2V). Diffusion models have revolutionized video generation, with AnimateDiff [4] and Video Diffusion Models [5] leading the way. Stable Video Diffusion [6] emphasized the importance of large, high-quality datasets for performance. Early models used U-Net backbones, but Sora [7] introduced the Diffusion Transformer (DiT) architecture [8], trained on extensive video corpora. Open-source models like CogVideoX [9], HunyuanVideo [10], and the WAN series [11] joined closed-source systems like Kling [12] and Veo 2 [13]. Most models share a common architecture: a 3D VAE compresses videos into spatiotemporal latents, and a DiT performs denoising. Data quality and scale remain crucial for model success, highlighting the importance of data curation and processing.
@@
Image-to-Video Generation (I2V). Early I2V systems extended T2V by conditioning on a single frame via latent concatenation or CLIP-based feature injection [14]. Later, cascaded designs like I2VGen-XL [15] and dual-injection frameworks like DynamiCrafter [16] introduced structured pipelines for better motion and appearance fidelity. Adapter-based approaches like LAMP [17] and I2V-Adapter [18] integrate cross-frame attention. Recent works on motion modeling and stable sampling include Motion-I2V [19] and FrameBridge [20]. Despite these advancements, I2V still faces challenges with curated datasets, limited long-range motion modeling, and the trade-off between appearance fidelity and motion realism.
@@
Text-to-Audio Generation (TTA). Recent advances in generative models have propelled text-to-audio generation. Models like Make-An-Audio [21] and AudioLDM [22] synthesize audio via iterative denoising of text-conditioned latent representations. Tango [23], Audio Flamingo [24], and others expand latent spaces and enhance cross-modal alignment. Stable Audio [25] employs hierarchical latent diffusion for high-fidelity output. AudioStory [26] introduces a unified generation framework, achieving long-form audio generation for the first time, while flow-matching techniques like VoiceBox [27] enable zero-shot style transfer. TangoFlux [28] optimizes text-audio alignment with CLAP-ranked preferences. Though these methods excel at semantic alignment, they are limited to short durations and lack flexibility for complex, evolving instructions, underscoring the need for TTA models that handle long, complex tasks.
@@
Audio-Video Joint Generation (T2AV). Pioneering efforts like MM-Diffusion [29] use coupled U-Net backbones, while DiT-based approaches dominate. AV-DiT [30] adapts pre-trained image DiTs with lightweight adapters, and UniForm [31] uses a unified single-tower architecture for audio-video tokens. Key challenges include precise spatio-temporal synchronization, addressed by methods like JavisDiT [1] (hierarchical prior), Ovi [3] (twin-backbone design), and SyncFlow [32] (dual-DiT with Rectified Flow Matching). Other research orchestrates unimodal experts, such as MMDisCo [33] and Universe-1 [2], which combine specialized models at the block level. Despite advances in architecture and data, most models focus on sound effects or music, leaving synchronized speech and video synthesis an underexplored challenge.
3 Klear
@@
Figure 1: Overview of Klear. The model takes four inputs: video, video-related text, audio-related text, and audio. Each input is individually encoded by respective encoders, then fed into the MM-DiT. The MM-DiT module outputs the latent variables of video and audio, which are then decoded separately into video and audio.
@@ 3.1 Preliminary
Conditional Flow-Matching. We employ flow matching as the denoising objective. The model needs to learn the velocity field that transforms pure noise $p_{0}=\mathcal{N}(0,\boldsymbol{I})$ to the underlying data distribution $p_{\text{data}}$. In practice, we perform linear interpolation $x_{t}=(1-t)x_{0}+tx_{1}$ to construct a distribution at timestamp $t$. Here, $x_{0}\sim p_{0}$ and $x_{1}\sim p_{\text{data}}$. Given the condition $c$, the model $\epsilon_{\theta}(\cdot)$ is trained to predict the target velocity, i.e., constantly as $u_{t}=x_{1}-x_{0}$:
@@
$\text{where}\quad t\sim\mathcal{U}(0,1),\ {x}_{0}\sim\mathcal{N}(0,\boldsymbol{I}),\ {x}_{1}\sim p_{\text{data}}.$
Latent Encoding. The model takes four inputs: video, video-related text, audio-related text, and audio, where video-related text represents the video caption and audio-related text represents the audio caption and speech text. Video is encoded by the 3d casual visual encoder from CogVideoX [9], We use Qwen3-8B Embedding [34] as the encoder for audio and video captions.
@@ 3.2 Single Tower with Full Attention
Single Tower DiT. To ensure a thorough audio-video fusion, we employ a single-tower architecture. As shown in Fig. 1, following Stable Diffusion 3 [35], we employ Multimodal Diffusion (MMDiT) to take the sequences of all modalities as input and perform full attention. Specifically, there are four inputs, i.e., video, video-related text, audio-related text, and audio. Each type of input is individually encoded into latents with respective encoders, then fed into the MM-DiT. The MM-DiT module outputs the latent variables of video and audio in two streams, which are then decoded separately to perform video and audio generation.
@@
Omni-Full Attention. Previous works may employ separated spatial and temporal attention to reduce computational complexity, like UniForm [31]. However, as in CogVideoX [9], this separate attention mechanism requires extensive implicit information transmission, significantly increasing the learning complexity. Other works tailor two transformer towers for audio and video generation separately, e.g., AV-DiT, SyncFlow, JavisDiT, TAVGBench.However, they often adopt a multi-stage training approach, which is complex and resource-intensive. The two towers must first be pretrained separately, then finetuned together, increasing training time and resource consumption. To achieve more efficient training and more effective modality fusion, we employ the 3D text-video-audio hybrid full attention mechanism. As shown in Fig. 1, within the MM-DiT module, the hidden states of video, video-related text, audio-related text, and audio are first scaled and normalized, then concatenated together for the attention calculation.
@@
$Attn(Q,K,V)=Softmax(\frac{QK^{\top}}{\sqrt{d_{k}}})V,$ (6)
@@ 3.3 Multi-Task Progressive Training Strategy
Random Modality Masking. To learn generalizable and robust audio-visual representations for joint generation, we train the generative model with a broad spectrum of tasks. As a result, we propose to selectively adjust the mask of query and key for audio and video modalities. If we restrict the query and key to video embedding and video caption embedding, the model degenerates to a T2V model. Similarly, limiting the query and key to audio embedding and audio text embedding results in a T2A model. In this way, the model could not only handle joint generation, but also maintain the abilities of single-modality generation. Considering the scarcity of high-quality audio-video paired data, our method offers an alternative for training the T2VA model. We first pre-train Klear on T2V and T2A tasks, and then finetune our model on audio-video paired data to finally construct a T2VA model. The learning objectives for audio and video generation are in Eq. (7) and Eq. (8):
@@
where $Mask_{a}$ is used to extract the audio token from the combined noise representation and $Mask_{v}$ is used to extract the vision tokens. In summary, $\mathcal{L}_{T2A}$ and $\mathcal{L}_{T2V}$ denote the single-modality tasks of T2A and T2V. To learn the generalizable and robust world knowledge of audio-visual correlation, we also incorporate several tasks of T2AV, I2V and I2AV. Consequently, the overall multi-task learning objective is as follows:
@@ 5 Experiments
In this section, we first present the experimental setup and implementation details (Sec. 5.1), then compare Klear with diverse baselines across multiple tasks (Sec. 5.2), complemented by qualitative results (Sec. 5.3). We further conduct ablations on the unified single-tower architecture, multi-task versus progressive training, and the role of 3D RoPE for native FPS (Sec. 5.4), collectively validating the effectiveness of our approach.
@@ 5.1 Experimental Setup
Baseline Methods. We select two canonical types of methods, including (1) Cascaded generation: these methods typically employ sequential T2V+V2A or T2A+A2V. Here, we employ AudioLDM2 [37] to perform the T2A task, while OpenSora [38] for T2V. (2) Joint generation. Recent works like JavisDiT [1], UniVerse-1 [2] and Ovi [3] employ a dual-tower architecture with dedicated interaction layers between them.
Evaluation Metrics. To evaluate the model’s generation capabilities, we test T2AV tasks on video, audio, and audio-video consistency. Following Universe-1, we use Verse-Bench for T2AV tasks. For video quality, we report the Motion Score (MS) based on RAFT [39] optical flow for dynamic realism and the Aesthetic Score (AS), a composite metric from MANIQA [40], aesthetic-predictor-v2-5 [41], and Musiq [42] for visual fidelity. Identity preservation is measured by ID Consistency (ID), calculated via DINOV3 [43] feature similarity. For audio quality, we use Fréchet Distance (FD) and KL Divergence on mel-spectrograms from PANNs [44]. Semantic alignment is evaluated using the CLAP score [45]. For synchronization, we report AV-A (Audio-Video Alignment) distance from Synchformer [46] and SyncNet Confidence (SNC) score [47] for lip sync. Global cross-modal alignment is measured by ImageBind (IB).
@@
Klear (Ours) Unified T2AV 0.48 0.51 0.59 1.36 1.06 0.232 0.028 0.65 6.79 0.316

### v2

# Unified Multi-Task Audio-Video Joint Generation

@@ Abstract
Audio–video joint generation has progressed rapidly, yet substantial challenges still remain. Non-commercial approaches still suffer audio-visual asynchrony, poor lip–speech alignment, and unimodal degradation, which can be stemmed from weak audio–visual correspondence modeling, limited generalization, and scarce high-quality dense-caption data. To address these issues, we introduce Apollo and delve into three axes—model architecture, training strategy, and data curation. Architecturally, we adopt a single-tower design with unified DiT blocks and an Omni-Full Attention mechanism, achieving tight audio–visual alignment and strong scalability. Training-wise, we adopt a progressive multitask regime—random modality masking to joint optimization across tasks, and a multistage curriculum, yielding robust representations, strengthening A-V aligned world knowledge, and preventing unimodal collapse. For datasets, we present the first large-scale audio–video dataset with dense captions, and introduce a novel automated data-construction pipeline which annotates and filters millions of diverse, high-quality, strictly aligned audio–video–caption triplets. Building on this, Apollo scales to large datasets, delivering high-fidelity, semantically and temporally aligned, instruction-following generation in both joint and unimodal settings while generalizing robustly to out-of-distribution scenarios. Across tasks, it substantially outperforms prior methods by a large margin and achieves performance comparable to Veo 3, offering a unified, scalable path toward next-generation audio–video synthesis.
Figure 1: We propose Apollo, a unified audio–video generation framework which delivers high fidelity, strong semantic and temporal alignment, and reliable instruction following in both joint and unimodal settings, with robust OOD generalization. Across tasks (T2AV/TI2AV/TI2V/T2V/T2A), it attains performance comparable to Veo-3 among open-source models.
@@ 1 Introduction
In ancient Greek mythology, Apollo presides over music, poetry, and prophecy, embodying artistic creation; he was later imbued with solar iconography and depicted as driving the sun chariot to illuminate the world. While for human perception and narrative, “illumination” arises not only from visible light and shadow but also from audible sensation: vision furnishes structure and space, while hearing shapes rhythm and affect, and their interplay yields a fully immersive experience. Accordingly, vision and auditory are complementary and jointly indispensable modalities for depicting the real world. Consequently, audio–video joint generation has emerged as one of the most prominent trends in generative AI. Commercial systems such as Veo 3 and Sora 2 have achieved impressive performance with strong semantic alignment, while diverse open-source models with varying architectures are also emerging. Nevertheless, research on joint audio–visual generation remains nascent, and current models—even some commercial systems—still exhibit audio–visual asynchrony, lip–speech mismatches, and degradation in unimodal quality. We attribute several factors: (1) Architecture. Most T2AV models like JavisDiT (Liu et al., 2025), UniVerse-1 (Wang et al., 2025a) and Ovi (Low et al., 2025) employ a single-tower architecture with a cross-attention module, resulting in limited audio-video interaction and alignment. (2) Training strategies and data. Currently, mainstream methods perform single-task training, which might bring about biased representations and struggle to exploit underlying audio-video correlation and world knowledge.
@@
To address these issues, we propose Apollo, which introduces coordinated improvements at the architectural, learning-strategy, and data levels. Specifically, on the architectural side, we adopt a single-tower backbone with unified DiT blocks. Each block integrates an Omni-Full Attention module that jointly attends to four streams—audio, audio captions, video, and video captions—thereby facilitating cross-modal fusion and interaction, achieving tight A-V alignment and stronger coupling to textual conditions, which offers a higher scaling ceiling. For training, we design a progressive multitask training regime: random modality masking sustains joint optimization over T2AV/TI2AV/TI2V/T2V/T2A. A performance-adaptive pretrain–post train curriculum tunes data mixtures and quality, yielding robust, generalizable representations that exploit A/V correlations and world knowledge. From the data construction, we introduce an automated annotation pipeline that enables efficient model scaling. Building on these three insights, our model achieves strong performance in both joint and unimodal generation, delivering high fidelity, tight audio–visual alignment, natural outputs, and robust instruction following with favorable scaling behavior. On the Verse-Bench, it surpasses prior methods by a large margin and generalizes well to out-of-distribution (OOD) scenarios. Our main contributions are summarized as follows:
  • We introduce Apollo, a unified framework for multi-task audio–video joint generation that, to our knowledge, is the first model to achieve performance comparable to Veo 3. Apollo effectively resolves semantic and temporal audio–visual misalignment while delivering high-fidelity generation.
@@
  • Extensive experiments demonstrate that Apollo consistently excels in both unimodal and joint audio–video generation, outperforming prior state-of-the-art on the unimodal benchmark and the AV joint-generation benchmark consistently.
@@ 2 Related Works
Text-to-Video Generation (T2V). Diffusion models have revolutionized video generation, with AnimateDiff (Guo et al., 2023) and Video Diffusion Models (Ho et al., 2022) leading the way. Stable Video Diffusion (Blattmann et al., 2023) emphasized the importance of large, high-quality datasets for performance. Early models used U-Net backbones, but Sora (Liu et al., 2024d) introduced the Diffusion Transformer (DiT) architecture (Peebles & Xie, 2023), trained on extensive video corpora. Open-source models like CogVideoX (Yang et al., 2024), HunyuanVideo (Kong et al., 2024a), and the WAN series (Wan et al., 2025) joined closed-source systems like Kling (Wang et al., 2025b) and Veo 2 (Deepanway et al., 2023). Most models share a common architecture: a 3D VAE compresses videos into spatiotemporal latents, and a DiT performs denoising. Data quality and scale remain crucial for model success, highlighting the importance of data curation and processing.
@@
Image-to-Video Generation (I2V). Early I2V systems extended T2V by conditioning on a single frame via latent concatenation or CLIP-based feature injection (Liu et al., 2023b). Later, cascaded designs like I2VGen-XL (Zhang et al., 2023) and dual-injection frameworks like DynamiCrafter (Xing et al., 2024) introduced structured pipelines for better motion and appearance fidelity. Adapter-based approaches like LAMP (Notomi et al., 2015) and I2V-Adapter (Guo et al., 2024) integrate cross-frame attention. Recent works on motion modeling and stable sampling include Motion-I2V (Shi et al., 2024) and FrameBridge (Wang et al., 2024b). Despite these advancements, I2V still faces challenges with curated datasets, limited long-range motion modeling, and the trade-off between appearance fidelity and motion realism.
@@
Text-to-Audio Generation (TTA). Recent advances in generative models have propelled text-to-audio generation. Models like Make-An-Audio (Huang et al., 2023) and AudioLDM (Liu et al., 2023a) synthesize audio via iterative denoising of text-conditioned latent representations. Tango (Liu et al., 2024a), Audio Flamingo (Kong et al., 2024b), and others expand latent spaces and enhance cross-modal alignment. Stable Audio (Evans et al., 2025) employs hierarchical latent diffusion for high-fidelity output. AudioStory (Guo et al., 2025) introduces a unified generation framework, achieving long-form audio generation for the first time, while flow-matching techniques like VoiceBox (Le et al., 2023) enable zero-shot style transfer. TangoFlux (Hung et al., 2024) optimizes text-audio alignment with CLAP-ranked preferences. Though these methods excel at semantic alignment, they are limited to short durations and lack flexibility for complex, evolving instructions, underscoring the need for TTA models that handle long, complex tasks.
@@
Audio-Video Joint Generation (T2AV). Pioneering efforts like MM-Diffusion (Ruan et al., 2023) use coupled U-Net backbones, while DiT-based approaches dominate. AV-DiT (Wang et al., 2024a) adapts pre-trained image DiTs with lightweight adapters, and UniForm (Zhao et al., 2025) uses a unified single-tower architecture for audio-video tokens. Key challenges include precise spatio-temporal synchronization, addressed by methods like JavisDiT (Liu et al., 2025) (hierarchical prior), Ovi (Low et al., 2025) (twin-backbone design), and SyncFlow (Liu et al., 2024b) (dual-DiT with Rectified Flow Matching). Other research orchestrates unimodal experts, such as MMDisCo (Hayakawa et al., 2024) and Universe-1 (Wang et al., 2025a), which combine specialized models at the block level. Despite advances in architecture and data, most models focus on sound effects or music, leaving synchronized speech and video synthesis an underexplored challenge.
3 Apollo
@@
Figure 2: Overview of Apollo. The model takes four inputs: video, video-related text, audio-related text, and audio. Each input is individually encoded by respective encoders, then fed into the MM-DiT. The MM-DiT module outputs the latent variables of video and audio, which are then decoded separately into video and audio.
@@ 3.1 Preliminary
Conditional Flow-Matching. We employ flow matching as the denoising objective. The model needs to learn the velocity field that transforms pure noise $p_{0}=\mathcal{N}(0,\bm{I})$ to the underlying data distribution $p_{\text{data}}$. In practice, we perform linear interpolation $x_{t}=(1-t)x_{0}+tx_{1}$ to construct a distribution at timestamp $t$. Here, $x_{0}\sim p_{0}$ and $x_{1}\sim p_{\text{data}}$. Given the condition $c$, the model $\epsilon_{\theta}(\cdot)$ is trained to predict the target velocity, i.e., constantly as $u_{t}=x_{1}-x_{0}$:
@@
$\text{where}\quad t\sim\mathcal{U}(0,1),\ {x}_{0}\sim\mathcal{N}(0,\bm{I}),\ {x}_{1}\sim p_{\text{data}}.$
Latent Encoding. The model takes four inputs: video, video-related text, audio-related text, and audio, where video-related text represents the video caption and audio-related text represents the audio caption and speech text. Video is encoded by the 3d casual visual encoder from CogVideoX (Yang et al., 2024), We use Qwen3-8B Embedding (Zhang et al., 2025) as the encoder for audio and video captions.
@@ 3.2 Single Tower with Full Attention
Single Tower DiT. To ensure a thorough audio-video fusion, we employ a single-tower architecture. As shown in Fig. 2, following Stable Diffusion 3 (Esser et al., 2024), we employ Multimodal Diffusion (MMDiT) to take the sequences of all modalities as input and perform full attention. Specifically, there are four inputs, i.e., video, video-related text, audio-related text, and audio. Each type of input is individually encoded into latents with respective encoders, then fed into the MM-DiT. The MM-DiT module outputs the latent variables of video and audio in two streams, which are then decoded separately to perform video and audio generation.
@@
Omni-Full Attention. Previous works may employ separated spatial and temporal attention to reduce computational complexity, like UniForm (Zhao et al., 2025). However, as in CogVideoX (Yang et al., 2024), this separate attention mechanism requires extensive implicit information transmission, significantly increasing the learning complexity. Other works tailor two transformer towers for audio and video generation separately, e.g., AV-DiT, SyncFlow, JavisDiT, TAVGBench.However, they often adopt a multi-stage training approach, which is complex and resource-intensive. The two towers must first be pretrained separately, then finetuned together, increasing training time and resource consumption. To achieve more efficient training and more effective modality fusion, we employ the 3D text-video-audio hybrid full attention mechanism. As shown in Fig. 2, within the MM-DiT module, the hidden states of video, video-related text, audio-related text, and audio are first scaled and normalized, then concatenated together for the attention calculation.
@@
$\text{Attn}(Q,K,V)=\text{Softmax}(\frac{QK^{\top}}{\sqrt{d_{k}}})V,$ (6)
@@ 3.3 Multi-Task Progressive Training Strategy
Random Modality Masking. To learn generalizable and robust audio-visual representations for joint generation, we train the generative model with a broad spectrum of tasks. As a result, we propose to selectively adjust the mask of query and key for audio and video modalities. If we restrict the query and key to video embedding and video caption embedding, the model degenerates to a T2V model. Similarly, limiting the query and key to audio embedding and audio text embedding results in a T2A model. In this way, the model could not only handle joint generation, but also maintain the abilities of single-modality generation. Considering the scarcity of high-quality audio-video paired data, our method offers an alternative for training the T2VA model. We first pre-train Apollo on T2V and T2A tasks, and then finetune our model on audio-video paired data to finally construct a T2VA model. The learning objectives for audio and video generation are in Eq. equation 7 and Eq. equation 8:
@@
where $\text{Mask}_{a}$ is used to extract the audio token from the combined noise representation and $\text{Mask}_{v}$ is used to extract the vision tokens. In summary, $\mathcal{L}_{T2A}$ and $\mathcal{L}_{T2V}$ denote the single-modality tasks of T2A and T2V. To learn the generalizable and robust world knowledge of audio-visual correlation, we also incorporate several tasks of T2AV, I2V and I2AV. Consequently, the overall multi-task learning objective is as follows:
@@ 5 Experiments
In this section, we first present the experimental setup and implementation details (Sec. 5.1), then compare Apollo with diverse baselines across multiple tasks (Sec. 5.2), complemented by qualitative results (Sec. 5.3). We further conduct ablations on the unified single-tower architecture, multi-task versus progressive training, and the role of 3D RoPE for native FPS (Sec. 5.4), collectively validating the effectiveness of our approach.
@@ 5.1 Experimental Setup
Baseline Methods. We select two canonical types of methods, including (1) Cascaded generation: these methods typically employ sequential T2V+V2A or T2A+A2V. Here, we employ AudioLDM2 (Liu et al., 2024c) to perform the T2A task, while OpenSora (Zheng et al., 2024) for T2V. (2) Joint generation. Recent works like JavisDiT (Liu et al., 2025), UniVerse-1 (Wang et al., 2025a) and Ovi (Low et al., 2025) employ a dual-tower architecture with dedicated interaction layers between them.
Evaluation Metrics. To evaluate the model’s generation capabilities, we test T2AV tasks on video, audio, and audio-video consistency. Following Universe-1, we use Verse-Bench for T2AV tasks. For video quality, we report the Motion Score (MS) based on RAFT (Teed & Deng, 2020) optical flow for dynamic realism and the Aesthetic Score (AS), a composite metric from MANIQA (Yang et al., 2022), aesthetic-predictor-v2-5 (discus0434, 2024), and Musiq (Ke et al., 2021) for visual fidelity. Identity preservation is measured by ID Consistency (ID), calculated via DINOV3 (Siméoni et al., 2025) feature similarity. For audio quality, we use Fréchet Distance (FD) and KL Divergence on mel-spectrograms from PANNs (Kong et al., 2020). Semantic alignment is evaluated using the CLAP score (Wu et al., 2023). For synchronization, we report AV-A (Audio-Video Alignment) distance from Synchformer (Iashin et al., 2024) and SyncNet Confidence (SNC) score (Chung & Zisserman, 2016) for lip sync. Global cross-modal alignment is measured by ImageBind (IB).
@@
Apollo (Ours) Unified T2AV 0.48 0.51 0.59 1.36 1.06 0.232 0.028 0.65 6.79 0.316